# Krylov Complexity as a Probe for Chaos

### v1

@@ I Introduction
The present article aims to explore the chaotic nature of many-body quantum systems using Krylov complexity [^1] Throughout this paper, we will examine the Krylov complexity of states Balasubramanian et al. (2022); Alishahiha and Banerjee (2023). , as defined in the context of the Lanczos algorithm. While this method has traditionally been employed for numerical computations Viswanath and Müller (1994), in recent years there has been increased interest in applying the Krylov method to the study of quantum chaos Parker et al. (2019).
@@
The presence of this peak before saturation has recently been observed numerically in various models Huh et al. (2024); Camargo et al. (2024); Nandy et al. (2024); Bhattacharjee and Nandy (2024); Balasubramanian et al. (2024); Baggioli et al. (2024). It is important to note that the appearance of a peak prior to saturation is not solely determined by the system’s dynamics; it also depends on the initial state. For instance, it occurs for the initial state considered in Erdmenger et al. (2023). Nonetheless, it appears that the thermofield double state at infinite temperature provides a maximal universal bound on this peak Camargo et al. (2024).
@@ IV Explicit examples
In the chaotic limit of this Ising model (43), these states have been considered in Bañuls et al. (2011) to explore the nature of thermalization, in particular, whether the thermalization is strong or weak [^8] By making use of certain measures the nature of thermalization for different spin models has been studied in Sun et al. (2021a); Lin and Motrunich (2017); Chen et al. (2021); Bhattacharjee et al. (2022b); dos Prazeres and de Oliveira (2024); Alishahiha and Vasli (2024).. In the strong thermalization, the expectation value of a typical operator relaxes to the thermal value very fast, while for weak thermalization it strongly oscillates around the thermal value, though its time average attains the thermal value. It has been shown that although the initial state $|Y+\rangle$ exhibits strong thermalization, the initial states $|Z+\rangle$ and $|X+\rangle$ do not (see also Sun et al. (2021b)).
@@
It is straightforward to compute the Lanczos coefficients for these initial states from which one can compute complexity numerically999 Lanczos coefficients for the model under consideration have also been computed inNoh (2021); Español and Wisniacki (2023); Trigueros and Lin (2022); Scialchi et al. (2024); Bhattacharya et al. (2023, 2022) . by using (19). The numerical results for the chaotic ($h=0.5$) and the integrable ($h=0$) cases are depicted in figures 1 and 2, respectively.
@@
It is worth noting once again that the actual value of the complexity saturation does not provide a measure to probe whether the model is chaotic or integrable. As evident from the results presented in figures 1 and 2, the infinite time average of complexity for the chaotic system is not significantly different from that of the integrable models101010see Rabinovici et al. (2022a, b); Scialchi et al. (2024); Ballar Trigueros and Lin (2022); Español and Wisniacki (2023) for related discussions.
@@ V Discussions
However, it is important to note that the exact value of complexity at the saturation point and its detailed behaviour at large and intermediate times depend not only on the dynamics but also on the initial state used for computing the Krylov complexity. Specifically, when the initial state is a maximally entangled state, the Krylov complexity, while approaching the infinite time average from above, leads to a peak before ultimately saturating at that average value. The appearance of this peak is not universal or state-independent. In fact, for a generic initial state, the complexity may or may not exhibit a clear peak. Moreover, there are small fluctuations around the saturation value whose shapes also depend on the initial state. This confirms the postulates of Susskind (2016); Brown et al. (2017) for complexity saturation.
@@
Finally, it is worth noting that for a thermal initial state where $|c_{j}|^{2}=e^{-\beta E_{j}}/Z[\beta]$, the Krylov complexity can be expressed in a form which is reminiscent of the spectral complexity introduced in Iliesiu et al. (2022). Therefore, we would expect that the spectral complexity exhibits the same late-time behaviour as that of the Krylov complexity.

### v2

@@ I Introduction
The present article aims to explore the chaotic nature of many-body quantum systems using Krylov complexity [^1] Throughout this paper, we will examine the Krylov complexity of states Balasubramanian et al. (2022); Alishahiha and Banerjee (2023); Caputa et al. (2024). , as defined in the context of the Lanczos algorithm. While this method has traditionally been employed for numerical computations Viswanath and Müller (1994), in recent years there has been increased interest in applying the Krylov method to the study of quantum chaos Parker et al. (2019) (for review see Nandy et al. (2024a))  [^2] For applications of Krylov method to entanglement generation and quantum metrology see e.g. Chu et al. (2024)..
@@
The presence of this peak before saturation has recently been observed numerically in various models Huh et al. (2024); Camargo et al. (2024a); Nandy et al. (2024b); Bhattacharjee and Nandy (2024); Balasubramanian et al. (2024); Baggioli et al. (2024). It is important to note that the appearance of a peak prior to saturation is not solely determined by the system’s dynamics; it also depends on the initial state. For instance, it occurs for the initial state considered in Erdmenger et al. (2023). Nonetheless, it appears that the thermofield double state at infinite temperature provides a maximal universal bound on this peak Camargo et al. (2024a).
@@ IV Explicit examples
In the chaotic limit of this Ising model (43), these states have been considered in Bañuls et al. (2011) to explore the nature of thermalization, in particular, whether the thermalization is strong or weak [^9] By making use of certain measures the nature of thermalization for different spin models has been studied in Sun et al. (2021a); Lin and Motrunich (2017); Chen et al. (2021); Bhattacharjee et al. (2022b); Nandy et al. (2024c); dos Prazeres and de Oliveira (2024); Alishahiha and Vasli (2024); Menzler and Jha (2024).. In the strong thermalization, the expectation value of a typical operator relaxes to the thermal value very fast, while for weak thermalization it strongly oscillates around the thermal value, though its time average attains the thermal value. It has been shown that although the initial state $|Y+\rangle$ exhibits strong thermalization, the initial states $|Z+\rangle$ and $|X+\rangle$ do not (see also Sun et al. (2021b)).
@@
It is straightforward to compute the Lanczos coefficients for these initial states from which one can compute complexity numerically101010 Lanczos coefficients for the model under consideration have also been computed inNoh (2021); Español and Wisniacki (2023); Trigueros and Lin (2022); Scialchi et al. (2024); Bhattacharya et al. (2023, 2022) . by using (19). The numerical results for the chaotic ($h=0.5$) and the integrable ($h=0$) cases are depicted in figures 1 and 2, respectively.
@@
It is worth noting once again that the actual value of the complexity saturation does not provide a measure to probe whether the model is chaotic or integrable. As evident from the results presented in figures 1 and 2, the infinite time average of complexity for the chaotic system is not significantly different from that of the integrable models111111see Rabinovici et al. (2022a, b); Scialchi et al. (2024); Ballar Trigueros and Lin (2022); Español and Wisniacki (2023) for related discussions.
@@ V Discussions
However, it is important to note that the exact value of complexity at the saturation point and its detailed behaviour at large and intermediate times depend not only on the dynamics but also on the initial state used for computing the Krylov complexity. Specifically, when the initial state is a maximally entangled state, the Krylov complexity, while approaching the infinite time average from above, leads to a peak before ultimately saturating at that average value. The appearance of this peak is not universal or state-independent. In fact, for a generic initial state, the complexity may or may not exhibit a clear peak. To clarify, for chaotic systems, the infinite-time average of complexity surpasses the thermal Krylov complexity. Consequently, the expression ${\rm Tr}\left((\rho_{{\rm DE}}-\rho_{{\rm th}}){\cal N}\right)$ is non-negative. Therefore, we observe a more pronounced peak for those states where this expression is near zero.
Moreover, there are small fluctuations around the saturation value whose shapes also depend on the initial state. This confirms the postulates of Susskind (2016); Brown et al. (2017) for complexity saturation.
@@
Finally, it is worth noting that for a thermal initial state where $|c_{j}|^{2}=e^{-\beta E_{j}}/Z[\beta]$, the Krylov complexity can be expressed in a form which is reminiscent of the spectral complexity introduced in Iliesiu et al. (2022) (see also Camargo et al. (2024b)). Therefore, we would expect that the spectral complexity exhibits the same late-time behaviour as that of the Krylov complexity.